If you invest $5,000 annually for the next 20 years and earn 9% per year (20 deposits, annuity, AER rate) . .

A. How much will you have at the end of the 20th year (T20), if the first deposit is 1 year from now (at T1) ?
B. How much will you have at the end of the 20th year, if the first deposit is made immediately (at T0) ?
Answer #1

Future Value if first deposit is made 1 year from now=Annual amount/rate*((1+rate)^t-1)

Future Value if first deposit is made now=Annual amount/rate*((1+rate)^t-1)*(1+rate)

1.
=5000/9%*(1.09^20-1)=255800.598209905

2.
=5000/9%*(1.09^20-1)*1.09=278822.652048797